One number is 6 less than another number. If the sum of the two numbers is 150, find the numbers
Answer by stanbon(75887)

One number is 6 less than another number. If the sum of the two numbers is 150, find the numbers
--------------------
x = y-6
x + y = 150
----
Rearrange:
x - y = -6
x + y = 150
----
2x = 144
x = 72
---
Solve for "y":
x + y = 150
72 + y = 150
y = 78
======================
Cheers,
Stan H.